# CardioAI: A Multimodal AI-based System to Support Symptom Monitoring and Risk Detection of Cancer Treatment-Induced Cardiotoxicity

### v2

@@ Abstract.
Despite recent advances in cancer treatments that prolong patients’ lives, treatment-induced cardiotoxicity remains one severe side effect. The clinical decision-making of cardiotoxicity is challenging, as non-clinical symptoms can be missed until life-threatening events occur at a later stage, and clinicians already have high workload centered on the treatment, not the side effect. Our project starts with a participatory design study with 11 clinicians to understand their practices and needs; then we build a multimodal AI system, CardioAI, that integrates wearables and LLM-powered voice assistants to monitor multimodal non-clinical symptoms. Also, the system includes an explainable risk prediction module that can generate cardiotoxicity risk scores and summaries as explanations to support clinicians’ decision-making. We conducted a heuristic evaluation with four clinical experts and found that they all believe CardioAI integrates well into their workflow, reduces their information overload, and enables them to make more informed decisions.
Cancer treatment-induced cardiotoxicity, Multimodal AI system, Large Language Models
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Human computer interaction (HCI)ccs: Applied computing Health informatics
@@ 1. Introduction
Figure 1. CardioAI is a multimodal AI-based system, designed to support clinical decision-making for cancer treatment-induced cardiotoxicity. CardioAI offers two main functionalities: (1) symptom monitoring: leveraging wearables and LLM-based voice assistants to collect physiological data and self-reported symptoms in non-clinical settings; (2) risk prediction: providing AI-generated risk scores for cardiotoxicity with feature importance and summary. The interface depicts five main modules in CardioAI: (A) Patient Information;(B) Daily Summary; (C) Wearable Sensor Data; (D) Explainable AI-generated Risk Score; (E) Conversation Log
@@
Cancer treatments, such as chemotherapy, have seen remarkable advancements over the past few decades, leading to significantly improved survival rates across various types of cancer (e.g., breast cancer, sarcoma) (Hu et al., 2024; van der Sluis et al., 2023; Wang et al., 2013; Raymond et al., 2011; Schmid et al., 2018; Borghaei et al., 2015; Hodi et al., 2010). For example, due to advancements in chemotherapy regimens, 5-year survival rates of breast cancer patients have improved from 82.7% to 91% (Wang et al., 2020). However, advanced cancer treatments still have severe adverse effects (e.g., liver and kidney damage), with treatment-induced cardiotoxicity standing out as one of the most severe complications (Truong et al., 2014; Shaikh and Shih, 2012; Florescu et al., 2013). Treatment-induced cardiotoxicity refers to the various heart damage caused by cancer treatments, which can lead to minor or severe consequences such as heart failure, arrhythmia, and other cardiovascular diseases (Alvi et al., 2019; Baptiste et al., 2019). To put this into the breast cancer treatment context, cardiotoxicity has emerged as the leading cause of morbidity and mortality in long-term follow-up (Henry et al., 2018).
Recent studies have investigated early decision-making of treatment-induced cardiotoxicity (Curigliano et al., 2016; Conway et al., 2015). However, this is challenging due to the following four reasons. Firstly, the symptoms of cardiotoxicity are often subtle in the early stages, making it hard to recognize the onset. When more pronounced symptoms (e.g., edema) appear, the condition may have already progressed to an irreversible state (Kalam and Marwick, 2013; Lima et al., 2022). Secondly, symptoms may occur infrequently, and with cancer patients typically visiting their clinicians only every few weeks (Henry et al., 2008), symptoms that occur outside of clinical settings may go unnoticed (Rabinowitz and Peirson, 2006). Thirdly, when patients experience discomfort at home, the primary way clinicians learn about this is through  patient self-reporting, as there are no existing methods for automatic symptom tracking. But patients may underreport or misreport symptoms due to forgetfulness, misunderstanding, or downplaying their discomfort (Cardinale et al., 2000, 2006; Pai and Nahata, 2000). In addition, clinicians are often overwhelmed with the primary task of managing and optimizing cancer treatments. This ongoing high workload and pressure leave them with limited time and capacity to process comprehensive information that would allow early detection and decision-making regarding cardiotoxicity (Ewer and Ewer, 2010; Curigliano et al., 2016). Due to these factors, decision-making for cancer treatment-induced cardiotoxicity is often delayed, which can result in missed opportunities for early symptom detection and proactive risk prediction, which are critical for timely intervention and improved patient outcomes (Wu et al., 2024).
@@
To address these challenges, we aim to develop an AI-based system to support clinicians’ early decision-making of cancer treatment-induced cardiotoxicity. Two groups of literature are particularly inspiring: Remote Patient Monitoring (RPM) (Calambur et al., 2024; Nguyen et al., 2023; Zhang et al., 2023) and AI-based Clinical Decision Support Sytems (AI-CDSS) (Wang et al., 2023). RPM uses technologies (e.g., wearable devices or home sensors) to continuously collect patient health data (e.g., blood pressure) outside of clinical settings (Catalyst, 2018; of Health et al., 2008), which could be a promising solution to cope with the aforementioned challenges of monitoring non-clinical subtle early symptom data. Furthermore, recent HCI research has shown the potential of using LLM-based voice assistants (LLM-VA) as verbal communication interfaces to support patients to self-report symptoms (Mahmood et al., 2023; Zhou et al., 2024; Yang et al., 2024a; Li et al., 2024), which can be a powerful complement to traditional RPM for our user case. With all the multimodal data continuously collected via RPM and LLM-VA, it may pose an additional cognitive demand on the already high-workload clinician decision makers, that is where AI-CDSS may help by automatically organizing and analyzing the raw data, and presenting only a high-level risk prediction to the clinicians (Chiang et al., 2023). The challenge is how to design a successful AI-CDSS that is explainable and trustworthy to the clinicians and can integrate into their existing workflow. Some existing HCI literature (Plana et al., 2014; Zhang et al., 2024; Albahri et al., 2023) exemplified the best practices such as following the human-centered AI (HCAI) principles in design (Amershi et al., 2019) and engaging stakeholder through the entire design process, which inspires our methodology.
@@
In this paper, we propose a first-of-its-kind multimodal AI-based system that combines the strengths of RPM (continuous monitoring of multimodal symptom data through wearable devices and LLM-VA) and explainable AI-CDSS (offer a model-based predictive risk score and explanation generation). The design process also follows the principles of human-centered AI design (Amershi et al., 2019). In particular, our work aims to answer two research questions:
@@
  • RQ1: What challenges do clinicians face in early decision-making for cancer treatment-induced cardiotoxicity, and what technological design needs do they have?
@@
  • RQ2: How will clinicians perceive and interact with our initial design of the AI-based multimodal system for early decision-making in cancer treatment-induced cardiotoxicity?
From RQ1, we aim to gather insights about the specific challenges and information technologies clinicians encounter in today’s clinical decision-making workflow, from which we aim to digest the design requirements for future technologies to address these challenges. To answer RQ1, we conduct participatory design (PD) sessions with 11 clinicians to solicit their needs and engage them in the co-design process with our initial design sketch 3. We then develop an interactive prototype, CardioAI, with two functionalities: (1) symptoms monitoring: vital signs collected by wearables and qualitative symptoms self-reported via LLM-VA; (2) risk prediction: AI-generated predictive risk scores for cardiotoxicity and accompanying explanations such as summary and feature importance. Lastly, to answer RQ2, we conducted a heuristic evaluation with four clinicians to collect their feedback on the usability and functionality of our system. Data are collected from cognitive walkthroughs and think-aloud during their interaction with the system and from post-study interviews and usability surveys. The results suggest that CardioAI is perceived to be useful in collecting and providing additional patient health data, and can reduce clinicians’ workload for decision-making.
In summary, our contributions are as follow:
@@
  • Following the human-centered design principle and practices, we co-designed with stakeholders and developed a multimodal AI-based system prototype that integrates continuous monitoring and AI predictive models to support early decision-making of cancer treatment-induced cardiotoxicity.
@@
  • Through an evaluation of our prototype, we identified key design considerations for future implementations of multimodal AI-based systems in clinical settings.
@@ 2. Related Work
We conduct a comprehensive review of existing challenges associated with cancer treatment-induced cardiotoxicity, various technologies used for remote monitoring of cancer patients, as well as AI-based clinical decision support systems(AI-CDSS) in this context.
@@ 2.1. Challenges in Cancer Treatment-Induced Cardiotoxicity
However, it is difficult for healthcare providers to reliably and effectively detect early signals of treatment-induced cardiotoxicity for timely intervention (Cardinale et al., 2010; Kalam and Marwick, 2013). One challenge is that initial cardiac symptoms caused by cardiotoxicity could be very mild, leading to belated self-report from patients (Wittayanukorn et al., 2017; Mascolo et al., 2021; Kamphuis et al., 2019), which could hinder early diagnosis of cardiotoxicity by clinicians during and after cancer treatment (Kalam and Marwick, 2013). In addition, treatment-induced cardiotoxicity can manifest at varying times: during treatment, shortly after, or even years after cancer treatment completion (Pai and Nahata, 2000). This variability in onset makes it challenging to monitor and detect, particularly when symptoms arise outside the hospital setting (Sheppard et al., 2013). Primary care physicians for cancer patients are generally burdened with the critical tasks of treating cancer patients in the hospital (Ewer and Ewer, 2010; Curigliano et al., 2016), with limited bandwidth to provide clinical support to cancer survivors after cancer treatment has been completed (Wu et al., 2024). It is worth noting that prior work discovers a lack of standardized guidelines for clinicians to detect treatment-induced cardiotoxicity early with effective interventions (Talty et al., 2024; Moustafa et al., 2023); thus, novel approaches for effective and efficient treatment-induced cardiotoxicity monitoring and diagnosis are of high demand (Gharib and Burnett, 2002).
@@ 2.2. Remote patient monitoring for Cancer Care
Remote patient monitoring (RPM) involves leveraging digital technologies to collect health data from patients outside clinical settings and transmit this information to healthcare providers for evaluation  (Catalyst, 2018; of Health et al., 2008). In cancer care, RPM has greatly improved clinicians’ ability to monitor and manage patient health, providing a cost-effective solution (Vegesna et al., 2017; Kofoed et al., 2012). Such improvement is driven by the integration of telehealth systems (Heijmans et al., 2019). Telehealth systems, including the use of technology-driven platforms to provide different types of health-related information and services (Mechanic et al., 2017), have facilitated cancer patients’ access to healthcare consultations and personalized coaching, boosting patient engagement and adherence to treatment protocols (Larson et al., 2018; Cox et al., 2017; Chen et al., 2018). Patients can self-administer their telehealth systems and data can be recorded and transmitted to clinicians remotely using a mobile phone (Doyle-Lindrud, 2020; Grewal et al., 2020). For instance, wearable telehealth systems (Dias and Paulo Silva Cunha, 2018; Shaik et al., 2022) are increasingly adopted in clinical settings to provide continuous real-time monitoring of patients’ physiological data, offering more accuracy than self-reported symptoms (Hardcastle et al., 2020; Gresham et al., 2018; Baig et al., 2017). For example, fitness trackers and smartwatches can monitor vital signs that are critical for the early detection of cardiotoxicity, such as heart rate, physical activity levels, sleep patterns, oxygen saturation, and respiratory rates, which can offer a thorough understanding of patients’ health status (Hardcastle et al., 2020; Tadas et al., 2023).
@@ 2.3. Design of AI-Supported Clinical Decision Support Systems
Recent advances in AI predictive modeling shed light on a broad avenue for supporting clinicians’ clinical decision-making with accessible and efficient AI assistance (Zhang et al., 2024). Such technical advancement, in turn, led to a growing research interest from HCI experts and healthcare practitioners in the design of AI-CDSS (Chiang et al., 2023; Chang et al., 2022; Lal and Cheng, 2023; Ahmed et al., 2024; Kwan et al., 2022; Yagi et al., 2024). AI-CDSS systems can greatly improve the efficiency of clinicians’ clinical workflow, where clinicians have to process a large amount of complicated electronic health records (EHRs) data to make informed clinical decisions (Ewer et al., 2023; Steinhubl et al., 2013, 2015; Free et al., 2013). However, traditional AI-CDSS approaches rely on clinical results for risk prediction or serve only general-purpose communication, while facing reliability, accuracy, or privacy concerns in AI applications in healthcare care (Chang et al., 2022; Lal and Cheng, 2023; Ahmed et al., 2024).

### v3

@@ Abstract.
Despite recent advances in cancer treatments that prolong patients’ lives, treatment-induced cardiotoxicity (i.e., the various heart damages caused by cancer treatments) emerges as one major side effect. The clinical decision-making process of cardiotoxicity is challenging, as early symptoms may happen in non-clinical settings and are too subtle to be noticed until life-threatening events occur at a later stage; clinicians already have a high workload focusing on the cancer treatment, no additional effort to spare on the cardiotoxicity side effect. Our project starts with a participatory design study with 11 clinicians to understand their decision-making practices and their feedback on an initial design of an AI-based decision-support system. Based on their feedback, we then propose a multimodal AI system, CardioAI, that can integrate wearables data and voice assistant data to model a patient’s cardiotoxicity risk to support clinicians’ decision-making. We conclude our paper with a small-scale heuristic evaluation with four experts and the discussion of future design considerations.
Human-AI collaboration, Cancer treatment-induced cardiotoxicity, Multimodal AI system, Large Language Models
journalyear: 2025copyright: ccconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3714272isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Human computer interaction (HCI)ccs: Applied computing Health informatics
Figure 1. CardioAI: a multimodal AI system to support clinicians for remote monitoring and risk detection of cancer patients’ cardiotoxicity risk. The UI has five modules: (A) Patient Information; (B) AI-generated Daily Summary; (C) Wearable Sensor Data; (D) AI-generated and Explainable Risk Score; (E) Conversation Log.
[…]
This figure depicts the main interface of the cardiotoxicity monitoring dashboard, designed to support clinicians by integrating patient data with AI predictions. The top left section, labeled A, displays the patient’s details, including name (Emily Johnson), age (52 y.o.), gender (Female), cancer type (Breast Cancer), cancer stage (IIA), and treatment type (Chemotherapy). Below this, the B section presents Self-Reported Symptoms, such as chest discomfort, as well as Wearable Sensor Data, showing metrics like heart rate (80.5 BPM average) and SpO2 levels (97% average). The Notes fields allow for additional information to be entered. In the C section, a graph shows trends in wearable sensor data, such as heart rate and respiration, over the past 24 hours. The D section to the right presents the Cardiotoxicity Risk Score, currently at 70%, and a Feature Importance breakdown, where chest discomfort contributes 50%, heart rate contributes 25%, and respiration contributes 15% to the overall risk prediction. The E section below includes a conversational log, where the chatbot asks the patient questions, and the patient responds about their current symptoms, such as feeling sleepy or reporting discomfort.
@@ 1. Introduction
Cancer treatments, such as chemotherapy, have seen remarkable advancements over the past few decades, leading to significantly improved survival rates across various types of cancer (e.g., breast cancer and sarcoma) (Hu et al., 2024; van der Sluis et al., 2023; Wang et al., 2013; Raymond et al., 2011; Schmid et al., 2018; Borghaei et al., 2015; Hodi et al., 2010). For example, thanks to advancements in chemotherapy regimens, 5-year survival rates of breast cancer patients have improved from 82.7% to 91% (Wang et al., 2020). However, advanced cancer treatments still have adverse effects (e.g., liver and kidney damage), with treatment-induced cardiotoxicity standing out as one of the most severe issues (Truong et al., 2014; Shaikh and Shih, 2012; Florescu et al., 2013). Treatment-induced cardiotoxicity refers to the various heart damage caused by cancer treatments, which can lead to minor or severe consequences such as heart failure, arrhythmia, and other cardiovascular diseases (Alvi et al., 2019; Baptiste et al., 2019). To put this into the breast cancer treatment context, cardiotoxicity has emerged as the leading cause of morbidity and mortality in long-term follow-up (Henry et al., 2018).
Recent studies have investigated early decision-making of treatment-induced cardiotoxicity (Curigliano et al., 2016; Conway et al., 2015). However, this is challenging due to the following four reasons. First, the initial symptoms of cardiotoxicity are often subtle in the early stages, and often, by the time clinicians are able to identify by obvious symptoms, cardiotoxicity has already caused irreversible damage (Kalam and Marwick, 2013; Lima et al., 2022)  (Challenge 1). Second, symptoms can manifest outside clinical settings, where patient-clinician interactions are significantly reduced, causing symptoms to go unnoticed or be diagnosed late (Henry et al., 2008) (Challenge 2). Third, the current healthcare system heavily relies on patient self-reporting for symptom tracking outside the hospital, but patients may falsely report and have poor compliance due to lack of health literacy and forgetfulness (Cardinale et al., 2000, 2006; Pai and Nahata, 2000) (Challenge 3). Lastly, clinicians face high workload and pressure, which limits their capacity to process massive amounts of clinical information in a short period of time necessary for early cardiotoxicity detection (Ewer and Ewer, 2010; Curigliano et al., 2016) (Challenge 4). While these challenges have been previously identified, their specific impact on clinicians’ decision-making processes remains underexplored. Without addressing these challenges, the identification of cancer treatment-induced cardiotoxicity is very difficult and often delayed, which could lead to belated disease intervention, and significantly impact overall patient outcome (Wu et al., 2024).
@@
To mitigate these challenges, we aim to develop an AI-based system to support clinicians’ early decision-making of cancer treatment-induced cardiotoxicity. Two groups of literature are particularly inspiring: Remote Patient Monitoring (RPM) (Calambur et al., 2024; Nguyen et al., 2023; Zhang et al., 2023) and AI-based Clinical Decision Support Systems (AI-CDSS) (Wang et al., 2023). RPM uses technologies (e.g., wearable devices or home sensors) to continuously collect patient health data (e.g., blood pressure) outside of clinical settings (Catalyst, 2018; of Health et al., 2008), which could be a promising solution to cope with the aforementioned challenges of monitoring non-clinical subtle early symptom data (Response to Challenge 1 & 2). Furthermore, recent HCI research has shown the potential of using LLM-based voice assistants (LLM-VA) as verbal communication interfaces to support patients to self-report symptoms (Mahmood et al., 2023; Zhou et al., 2024; Yang et al., 2024a; Li et al., 2024), which can be a powerful complement to traditional self-reporting approaches for our user case  (Response to Challenge 3). With all the multimodal data continuously collected via RPM and LLM-VA, it may pose an additional cognitive demand on the already high-workload clinician decision makers; that is where AI-CDSS may help by automatically organizing and analyzing the raw data, and presenting only a high-level risk prediction to the clinicians (Chiang et al., 2023)  (Response to Challenge 4).
@@
One key issue of designing a successful AI-CDSS is that the AI needs to be explainable and trustworthy to clinicians, and it should integrate into their existing workflow. Some existing HCI literature (Plana et al., 2014; Zhang et al., 2024; Albahri et al., 2023) exemplified the best practices such as following the human-centered AI (HCAI) principles in design (Amershi et al., 2019) and engaging stakeholders throughout the design process (Wang et al., 2023; Zhang et al., 2024), which inspired our methodology. In this paper, we propose a first-of-its-kind multimodal AI-based system that combines the strengths of RPM (continuous monitoring multimodal symptom data via wearables and a LLM-VA to address Challenge 1, 2, and 3) and explainable AI-CDSS (designed to mitigate Challenge 4 by offering a model-based predictive risk score and explanation generation). The design process also follows HCAI design principles (Amershi et al., 2019). In particular, our work aims to answer two research questions:
@@
  • RQ1: How do clinicians’ experiences provide deeper insights into the challenges of early decision-making for cancer treatment-induced cardiotoxicity, and what technological design needs do they have?
@@
  • RQ2: How will clinicians perceive and interact with our initial design of a multimodal AI-based system in early decision-making of cancer treatment-induced cardiotoxicity?
From RQ1, we aim to gather clinicians’ challenges and technological needs in today’s decision-making workflow for cancer treatment-induced cardiotoxicity. To answer RQ1, we conducted participatory design (PD) sessions with 11 clinicians, fostering in-depth discussions and actively involving them in the co-design process. During each session (Figure 2), we presented an initial design sketch to gather feedback and suggestions. Through this collaborative and iterative process, we developed our pilot system, CardioAI, which is detailed in Section 4. CardioAI offers two main functionalities: (1)  Symptom Monitoring: vital signs collected from wearables (Module C in Figure  1) and qualitative symptoms self-reported via an LLM-VA (Module E); (2)  Risk Prediction and Summarization: AI-generated predictive risk scores for cardiotoxicity paired with explanations as feature importance (Module D) and summarizations of the collected data (Module B).
Lastly, to address RQ2, we conducted a heuristic evaluation with four clinicians, including two cardiologists and two oncologists, to gather feedback on the usability and functionality of our pilot system. Data were collected through cognitive walkthroughs and think-aloud protocols during clinicians’ interactions with the system, along with post-study interviews and usability surveys. Clinicians highlighted the system’s ability to provide additional cardiotoxicity-specific patient data for monitoring and diagnosis, reduce cognitive workload, and support proactive decision-making within their existing workflows.
In summary, our contributions are as follows:
@@
  • Following the human-centered design principles and practices, we co-designed with clinicians a multimodal AI-based pilot system that integrates continuous symptom monitoring and explainable AI model to support early decision-making of cancer treatment-induced cardiotoxicity.
@@
  • Through a follow-up user evaluation of our prototype, we identified key design considerations for future implementations of multimodal AI-based systems in clinical settings.
@@ 2.1. Challenges in Cancer Treatment-Induced Cardiotoxicity
However, detecting early signals of treatment-induced cardiotoxicity for timely intervention is a significant challenge in the healthcare field (Cardinale et al., 2010; Kalam and Marwick, 2013). One major issue is that initial cardiac symptoms are often very mild, leading to delayed self-reporting from patients (Wittayanukorn et al., 2017; Mascolo et al., 2021; Kamphuis et al., 2019) (Challenge 1). When more pronounced symptoms (e.g., edema) appear, the condition may have already progressed to an irreversible state (Kalam and Marwick, 2013; Lima et al., 2022). This delay hinders clinicians’ ability to diagnose cardiotoxicity early during and after cancer treatment (Kalam and Marwick, 2013). Additionally, treatment-induced cardiotoxicity can manifest at varying times—during treatment, shortly after, or even years later (Pai and Nahata, 2000). This variability makes consistent monitoring challenging, especially when symptoms arise outside the hospital setting (Sheppard et al., 2013)  (Challenge 2). The reliance on patient self-reporting is another critical challenge. Patients may underreport or misreport symptoms due to forgetfulness, misunderstanding, or downplaying their discomfort (Cardinale et al., 2000, 2006; Pai and Nahata, 2000). This underreporting can lead to missed early signs of cardiotoxicity, delaying necessary interventions (Challenge 3).
Furthermore, clinicians are frequently tasked with the demanding responsibility of overseeing cancer treatments (Ewer and Ewer, 2010; Curigliano et al., 2016), leaving them with little capacity to rapidly gather and analyze extensive clinical data, which is an essential step for the early detection of cardiotoxicity (Wu et al., 2024) (Challenge 4). There is also a lack of standardized guidelines for clinicians to detect and intervene in treatment-induced cardiotoxicity effectively (Talty et al., 2024; Moustafa et al., 2023), which exacerbates the difficulty in early detection and consistent management across healthcare providers. Due to these compounded challenges, there is a high demand for novel approaches that offer effective and efficient monitoring and diagnosis of treatment-induced cardiotoxicity (Gharib and Burnett, 2002).
@@ 2.2. Remote patient monitoring for Cancer Care
Remote patient monitoring (RPM) involves leveraging digital technologies to collect health data from patients outside clinical settings and transmit this information to healthcare providers for evaluation  (Catalyst, 2018; of Health et al., 2008). In cancer care, RPM has greatly improved clinicians’ ability to monitor and manage patient health, providing a cost-effective solution (Vegesna et al., 2017; Kofoed et al., 2012). Such improvement is driven by the integration of telehealth systems (Heijmans et al., 2019). Telehealth systems, including the use of technology-driven platforms to provide different types of health-related information and services (Mechanic et al., 2017), have facilitated cancer patients’ access to healthcare consultations and personalized coaching, boosting patient engagement and adherence to treatment protocols (Larson et al., 2018; Cox et al., 2017; Chen et al., 2018). Patients can self-administer their telehealth systems, and data can be recorded and transmitted to clinicians remotely using a mobile phone (Doyle-Lindrud, 2020; Grewal et al., 2020). For instance, wearable telehealth systems (Dias and Paulo Silva Cunha, 2018; Shaik et al., 2022) are increasingly adopted in clinical settings to provide continuous real-time monitoring of patients’ physiological data, offering more accuracy than self-reported symptoms (Hardcastle et al., 2020; Gresham et al., 2018; Baig et al., 2017). For example, fitness trackers and smartwatches can monitor vital signs that are critical for the early detection of cardiotoxicity, such as heart rate, physical activity levels, sleep patterns, oxygen saturation, and respiratory rates, which can offer a thorough understanding of patients’ health status (Hardcastle et al., 2020; Tadas et al., 2023).
@@ 2.3. Design of AI-Supported Clinical Decision Support Systems
Recent advances in AI predictive modeling shed light on a broad avenue for supporting clinicians’ clinical decision-making with accessible and efficient AI assistance (Zhang et al., 2024). Such technical advancement, in turn, led to a growing research interest from HCI experts and healthcare practitioners in the design of AI-CDSS (Chiang et al., 2023; Chang et al., 2022; Lal and Cheng, 2023; Ahmed et al., 2024; Kwan et al., 2022; Yagi et al., 2024). AI-CDSS systems can greatly improve the efficiency of clinicians’ clinical workflow, where clinicians have to process a large amount of complicated electronic health records (EHRs) data to make informed clinical decisions (Ewer et al., 2023; Steinhubl et al., 2013, 2015; Free et al., 2013). However, traditional AI-CDSS approaches rely on clinical results for risk prediction or serve only general-purpose communication while facing reliability, accuracy, or privacy concerns in AI applications in healthcare care (Chang et al., 2022; Lal and Cheng, 2023; Ahmed et al., 2024).